# Hyperbolic recurrent neural network as the first type of non-Euclidean neural quantum state ansatz

### v1

@@ Abstract
In this work, we introduce the first type of non-Euclidean neural quantum state (NQS) ansatz, in the form of the hyperbolic GRU (a variant of recurrent neural networks (RNNs)), to be used in the Variational Monte Carlo method of approximating the ground state wavefunction for quantum many-body systems. In particular, we examine the performances of NQS ansatzes constructed from both conventional or Euclidean RNN/GRU and from hyperbolic GRU in the prototypical settings of the one- and two-dimensional transverse field Ising models (TFIM) of up to 100 spins and the one-dimensional Heisenberg $J_{1}J_{2}$ and $J_{1}J_{2}J_{3}$ systems of up 50 spins with varying values of $(J_{1},J_{2})$ and $(J_{1},J_{2},J_{3})$. By virtue of the fact that, for all of the experiments performed in this work, hyperbolic GRU can yield performances comparable to or better than Euclidean RNNs, which have been extensively studied in these settings in the literature, our work is a proof-of-concept for the viability of hyperbolic GRU as the first type of non-Euclidean NQS ansatz for quantum many-body systems. Furthermore, in settings where the Hamiltonian displays a clear hierarchical interaction structure, such as the 1D Heisenberg $J_{1}J_{2}$ $J_{1}J_{2}J_{3}$ systems with the first, second and even third nearest neighbor interactions, our results show that hyperbolic GRU definitively outperforms its Euclidean version in all instances. Additionally, in the 2D TFIM setting, 1D hyperbolic GRU NQS also outperform 1D Euclidean GRU NQS, due to the presence of a hierarchical interaction structure comprising the first and $N^{th}$ nearest neighbor interactions, introduced by the rearrangement of the 1D spin chain to mimic the 2D lattice. The fact that these results are reminiscent of the established ones from natural language processing where hyperbolic GRU almost always outperforms Euclidean RNNs when the training data exhibit a tree-like or hierarchical structure leads us to hypothesize that hyperbolic GRU NQS ansatz would likely outperform Euclidean RNN/GRU NQS ansatz in quantum spin systems that involve different degrees of nearest neighbor interactions. Finally, with this work, we hope to initiate future studies of other types of non-Euclidean neural quantum states beyond hyperbolic GRU.
@@ 5.1.2 Results
Figure 4: VMC experiments for 1D TFIM with $N=20,40,80$ and 100 spins (from top to bottom row, left to right): Comparisons of the performances of eRNN-50-s50, eGRU-50-s50, hypGRU-50-s50 - abbreviated as eRNN, eGRU and hGRU on the horizontal axis of each subfigure - during training and inference processes. In each subfigure, the mean energy from training and inference for each NQS ansatz is shown as a dot with error bar (representing the standard error).
@@ 5.2.2 Results
From Table 5, Table 6 and Fig.6, it is clear that 2D Euclidean RNN is the best performing ansatz in all four cases for both training and inference, since the architecture of the RNN part of this NQS was specially created for the 2D setting of TFIM. It is thus within our expectation that 2D RNN will outperform both 1D Euclidean and hyperbolic GRU. What we are really interested in, and would like to find out, however, is whether 1D hyperbolic GRU will outperform 1D Euclidean GRU. Regarding this aspect, we observe that in the training process, hyperbolic GRU outperform Euclidean GRU 3 out of 4 times when $(N_{x},N_{y})=(5,5),(7,7),(9,9)$, while in inference process, hyperbolic GRU outperform Euclidean GRU 4 out of 4 times (see Fig.6). While for $(N_{x},N_{y})=(5,5),(7,7)$, it is clear that hyperbolic GRU definitively outperform Euclidean GRU for in both training and inference, the situation is not clear-cut for $(N_{x},N_{y})=(8,8),(9,9)$. When $(N_{x},N_{y})=(9,9)$ Euclidean and hyperbolic GRU are almost comparable in performance during training, while for $N_{x},N_{y})=(8,8)$, their inference results are almost comparable, although for both cases, hyperbolic GRU still emerged as the slightly better variant. Also, from the converge curve for the mean energy, Fig.12, we note that the cyan curve corresponding to 1D hyperbolic GRU is always closer, with narrower range of fluctuations, to the DMRG line than the blue curve corresponding to 1D Euclidean GRU, even for the case of $(N_{x},N_{y})=(8,8)$, when the final saved 1D hyperbolic GRU model does not yield the result as good as that of 1D Euclidean GRU (see the first subfigure on the second row of Fig.12). While more experiments are needed, especially for larger $(N_{x},N_{y})$ sizes, from the results of these experiments, we see that 1D hyperbolic GRU is likely to outperform 1D Euclidean GRU.
Figure 6: Comparisons of the performances (during training and inference) of 2D Euclidean RNN and 1D Euclidean/hyperbolic GRU ansatzes for the 2D TFIM with different ($N_{x},N_{y}$) lattices - clockwise from top: $(N_{x},N_{y})=(5,5)$, (7,7), (9,9), (8,8). Each dot with error bar represents the mean $E$ value of an NQS ansatz. For all cases, 2D Euclidean RNN is the best ansatz. The second best ansatz is 1D hyperbolic GRU for all cases, except for the training result of $(N_{x},N_{y})=(8,8)$.
@@ 5.3.1 Experiment details
Table 7: The variants of Euclidean and hyperbolic GRU used as NQS ansatzes to run the VMC experiments involving the 1D $J_{1}J_{2}$ Hamiltonian Eq.(53) with $N=50$, $J_{1}=1$ and $J_{2}=0.0,0.2,0.5,0.8$. The second and last rows titled ‘Parameters’ list the numbers of parameters in the Euclidean and hyperbolic GRU networks.
@@ 5.3.2 Results
Figure 8: Comparisons of the performances of Euclidean and hyperbolic GRU NQS ansatzes from VMC runs of 1D $J_{1}J_{2}$ model with different $J_{2}$ values (clockwise from top: $J_{2}$ =0.0, 0.2, 0.8, 0.5) for $N=50$ spins. In each subfigure, the mean energy and standard error from training and inference for each NQS ansatz is shown as a dot with error bar. The specific variants of Euclidean/hyperbolic GRU used in each of the four cases are listed in Table 7. In this plot, we do not specify the exact variants, but label them simply as Euclidean GRU or hyperbolic GRU to emphasize the geometry of the ansatzes.
@@ 5.4.1 Experiment details
The NQS used are the complex Euclidean GRU and hyperbolic GRU ansatzes as in the 1D $J_{1}J_{2}$ case, with the RNN layer having the hidden dimension of $d_{h}=50$ (see Table 10 for the variant names and their number of parameters).
@@
Table 10: The variants of Euclidean and hyperbolic GRU used as NQS ansatzes to run the VMC experiments involving the 1D $J_{1}J_{2}J_{3}$ Hamiltonian Eq.(55) with $N=30$, $J_{1}=1$ and $(J_{2},J_{3})=(0.0,0.5),(0.2,0.2),(0.2,0.5),(0.5,0.2)$. The number of parameters of each NQS is listed in brackets under their names. Only one variant for each type of NQS ansatz is used in this set of VMC experiments.
Unlike the 1D $J_{1}J_{2}$ case where three different variants of hyperbolic GRU with hidden dimensions $d_{H}=65,70,75$ were used and the best performing variant chosen to compare against the Euclidean variant with $d_{h}=75$, in this case, we only use one variant each for either Euclidean or hyperbolic GRU, with the fixed hidden dimension of $d_{h}=50$, due to our limited computing resources. In each VMC experiment, each ansatz is trained for a fixed number of 280 epochs. Euclidean GRU typically takes around 3 hours to complete the training (on a laptop with 16Gb of RAM and no GPU), while hyperbolic GRU takes more than 24 hours to complete the training (on the same hardware). All other training details (optimizer settings, saving-of-best-model-only method) remain the same as in the 1D $J_{1}J_{2}$ case and will not be repeated here.
@@ 5.4.2 Results
The training and inference results are listed in Table 11 and Table 12, respectively. Fig.9 graphically shows the data in both these Tables and offers a comparison between the performances of the 1D Euclidean and 1D hyperbolic GRU. The convergence curve plots for the mean energy and energy variance recorded during the training process are shown in Fig.16 and Fig.17, respectively, in Appendix A.4.
@@
Table 11: VMC training results for the 1D $J_{1}J_{2}J_{3}$ system with the Hamiltonian Eq.(55) with $N=30$, $J_{1}=1$, and $(J_{2},J_{3})=(0.0,0.2),(0.2,0.2),(0.2,0.5),(0.5,0.2)$. These results are recorded directly from the training and correspond to the parameters of the best saved model. For each ansatz, the mean $E$ is listed first, together with the variance (in brackets). The best result at each $(J_{2},J_{3})$ case is noted in bold.
@@
Table 12: VMC inference results for the 1D $J_{1}J_{2}J_{3}$ system with the Hamiltonian Eq.(55) with $N=30$, $J_{1}=1$, and $(J_{2},J_{3})=(0.0,0.2),(0.2,0.2),(0.2,0.5),(0.5,0.2)$. For each ansatz, the mean $E$ is listed first, together with the variance (in brackets). The best result at each $(J_{2},J_{3})$ case is noted in bold.
@@
Figure 9: Comparisons of the performances of Euclidean and hyperbolic GRU NQS ansatzes from VMC runs of 1D $J_{1}J_{2}J_{3}$ model with different $(J_{2},J_{3})$ values - clockwise from top: (0.0, 0.5), (0.2, 0.2), (0.5, 0.2), (0.2, 0.5) - for $N=30$ spins. In each subfigure, the mean energy and standard error from training and inference for each NQS ansatz is shown as a dot with an error bar. As is the $J_{1}J_{2}$ case, we do not specify the exact variants (listed in Table 10) in this plot, but choosing instead to label them as ‘Euclidean GRU’ and ‘Hyperbolic GRU’ to emphasize the geometry of the NQS ansatz.
@@
From Table 11, Table 12 and Fig.9, we note that in both training and inference, hyperbolic GRU definitively outperformed Euclidean GRU (by a large margin) in all 4 out of 4 cases. The energy convergence curves (see Fig.16) corresponding to the hyperbolic GRU ansatz reach convergence at values well below that reached by the curves corresponding to the Euclidean GRU ansatz, while the variance convergence curves of hyperbolic GRU (see Fig.17) show smaller ranges of fluctuations than Euclidean GRU’s. These results are in agreement with the ones in the 1D $J_{1}J_{2}$ case in which hyperbolic GRU outperfomed its Euclidean version in all cases when $J_{2}\neq 0$, i.e. when the second neighbor interaction is nonvanishing and the Hamiltonian exhibits a hierarchy in the interaction structure comprised of the first and second neighbor interactions. In all four $J_{1}J_{2}J_{3}$ settings where there exist an even more pronounced hierarchy in the interaction structure of the Hamiltonian that comprises the first, second and third nearest neighbor interactions, the clear outperformance of the hyperbolic GRU ansatz reinforces the hypothesis that the hyperbolic geometry of this neural network plays a determining role in handling the type of Hamiltonians with hierarchical structures such as the 1D $J_{1}J_{2}$, $J_{1}J_{2}J_{3}$ systems. As already noted in the 1D $J_{1}J_{2}$ case, an interesting task would be to verify whether hyperbolic GRU carries on outperforming Euclidean GRU for the 2D $J_{1}J_{2}J_{3}$ models on different lattice geometries, but this will require the generalization of the 1D hyperbolic GRU network to a 2D version, similar to the generalization of the 1D Euclidean RNN cell to the 2D version in Eq.52 as done in [9].

### v2

@@ Abstract
In this work, we introduce the first type of non-Euclidean neural quantum state (NQS) ansatz, in the form of the hyperbolic GRU (a variant of recurrent neural networks (RNNs)), to be used in the Variational Monte Carlo method of approximating the ground state wavefunction for quantum many-body systems. In particular, we examine the performances of NQS ansatzes constructed from both conventional or Euclidean RNN/GRU and from hyperbolic GRU in the prototypical settings of the one- and two-dimensional transverse field Ising models (TFIM) and the one-dimensional Heisenberg $J_{1}J_{2}$ and $J_{1}J_{2}J_{3}$ systems with varying values of $(J_{1},J_{2})$ and $(J_{1},J_{2},J_{3})$. By virtue of the fact that, for all of the experiments performed in this work, hyperbolic GRU can yield performances comparable to or better than Euclidean RNNs, which have been extensively studied in these settings in the literature, our work is a proof-of-concept for the viability of hyperbolic GRU as the first type of non-Euclidean NQS ansatz for quantum many-body systems. Furthermore, in settings where the Hamiltonian displays a clear hierarchical interaction structure, such as the 1D Heisenberg $J_{1}J_{2}$ $J_{1}J_{2}J_{3}$ systems with the first, second and even third nearest neighbor interactions, our results show that hyperbolic GRU definitively outperforms its Euclidean version in almost all instances. Additionally, in the 2D TFIM setting, 1D hyperbolic GRU NQS also outperform 1D Euclidean GRU NQS, due to the presence of a hierarchical interaction structure comprising the first and $N^{th}$ nearest neighbor interactions, introduced by the rearrangement of the 1D spin chain to mimic the 2D lattice. The fact that these results are reminiscent of the established ones from natural language processing where hyperbolic GRU almost always outperforms Euclidean RNNs when the training data exhibit a tree-like or hierarchical structure leads us to hypothesize that hyperbolic GRU NQS ansatz would likely outperform Euclidean RNN/GRU NQS ansatz in quantum spin systems that involve different degrees of nearest neighbor interactions. Finally, with this work, we hope to initiate future studies of other types of non-Euclidean neural quantum states beyond hyperbolic GRU.
@@ 5.1.2 Results
Figure 4: VMC experiments for 1D TFIM with $N=20,40,80$ and 100 spins (from top to bottom row, left to right): Comparisons of the performances of the variants eRNN-50-s50, eGRU-50-s50, hypGRU-50-s50 (listed in Table 1) - abbreviated as eRNN, eGRU and hGRU on the horizontal axis of each subfigure - during training and inference processes. In each subfigure, the mean energy from training and inference for each NQS ansatz is shown as a dot with error bar (representing the standard error).
@@ 5.2.2 Results
From Table 5, Table 6 and Fig.6, it is clear that 2D Euclidean RNN is the best performing ansatz in all four cases for both training and inference, since the architecture of the RNN part of this NQS was specially created for the 2D setting of TFIM. It is thus within our expectation that 2D RNN will outperform both 1D Euclidean and hyperbolic GRU. What we are really interested in, and would like to find out, however, is whether 1D hyperbolic GRU will outperform 1D Euclidean GRU. Regarding this aspect, we observe that in the training process, hyperbolic GRU outperform Euclidean GRU 3 out of 4 times when $(N_{x},N_{y})=(5,5),(7,7),(9,9)$, while in inference process, hyperbolic GRU outperform Euclidean GRU 4 out of 4 times (see Fig.6). While for $(N_{x},N_{y})=(5,5),(7,7)$, it is clear that hyperbolic GRU definitively outperform Euclidean GRU for in both training and inference, the situation is not clear-cut for $(N_{x},N_{y})=(8,8),(9,9)$. When $(N_{x},N_{y})=(9,9)$ Euclidean and hyperbolic GRU are almost comparable in performance during training, while for $N_{x},N_{y})=(8,8)$, their inference results are almost comparable, although for both cases, hyperbolic GRU still emerged as the slightly better variant. Also, from the converge curve for the mean energy, Fig.13, we note that the cyan curve corresponding to 1D hyperbolic GRU is always closer, with narrower range of fluctuations, to the DMRG line than the blue curve corresponding to 1D Euclidean GRU, even for the case of $(N_{x},N_{y})=(8,8)$, when the final saved 1D hyperbolic GRU model does not yield the result as good as that of 1D Euclidean GRU (see the first subfigure on the second row of Fig.13). While more experiments are needed, especially for larger $(N_{x},N_{y})$ sizes, from the results of these experiments, it can be seen that 1D hyperbolic GRU is likely to outperform 1D Euclidean GRU.
Figure 6: Comparisons of the performances (during training and inference) of 2D Euclidean RNN and 1D Euclidean/hyperbolic GRU ansatzes listed in Table 4 for the 2D TFIM with different ($N_{x},N_{y}$) lattices - clockwise from top: $(N_{x},N_{y})=(5,5)$, (7,7), (9,9), (8,8). Each dot with error bar represents the mean $E$ value of an NQS ansatz. For all cases, 2D Euclidean RNN is the best ansatz. The second best ansatz is 1D hyperbolic GRU for all cases, except for the training result of $(N_{x},N_{y})=(8,8)$.
@@ 5.3.1 Experiment details
Table 7: The variants of Euclidean and hyperbolic GRU used as NQS ansatzes to run the VMC experiments involving the 1D $J_{1}J_{2}$ Hamiltonian Eq.(53) with $N=50$, $J_{1}=1$ and $J_{2}=0.0,0.2,0.5,0.8$. While three variants of hyperbolic GRU with $d_{h}=60,70,75$ were used in the experiments, we only list the best performing variant in this Table. The second and last rows titled ‘Parameters’ list the numbers of parameters in the Euclidean and hyperbolic GRU networks.
@@ 5.3.2 Results
Figure 8: Comparisons of the performances of Euclidean and hyperbolic GRU NQS ansatzes listed in Table 7 from VMC runs of 1D $J_{1}J_{2}$ model with different $J_{2}$ values (clockwise from top: $J_{2}$ =0.0, 0.2, 0.8, 0.5) for $N=50$ spins. In each subfigure, the mean energy and standard error from training and inference for each NQS ansatz is shown as a dot with error bar. The specific variants of Euclidean/hyperbolic GRU used in each of the four cases are listed in Table 7. In this plot, we do not specify the exact variants, but label them simply as Euclidean GRU or hyperbolic GRU to emphasize the geometry of the ansatzes.
@@ 5.4.1 Experiment details
The NQS used are the complex Euclidean GRU and hyperbolic GRU ansatzes as in the 1D $J_{1}J_{2}$ case. We will perform two different sets of VMC experiments.
  • In the first set of experiments, all four ($J_{2},J_{3}$) cases listed above are considered, withnthe Euclidean/hyperbolic GRU NQS ansatzes used having the same RNN hidden dimension of $d_{h}=50$ (see Table 10 for the variant names and their numbers of parameters).
@@
    Table 10: The variants of Euclidean and hyperbolic GRU used as NQS ansatzes to run the first set of VMC experiments involving the 1D $J_{1}J_{2}J_{3}$ Hamiltonian Eq.(55) with $N=30$, $J_{1}=1$ and $(J_{2},J_{3})=(0.0,0.5),(0.2,0.2),(0.2,0.5),(0.5,0.2)$. The number of parameters of each NQS is listed in brackets under their names. Only one variant for each type of NQS ansatz is used in this set of VMC experiments.
  • In the second set of experiments, we only considered $(J_{2},J_{3})$ = (0.2, 0.5), (0.5, 0.2), with the Euclidean GRU ansatz having the RNN hidden dimension $d_{h}=60$ and the hyperbolic GRU ansatz having $d_{h}=55$ (see Table 11).
    ($J_{2},J_{3})$ Euclidean GRU Hyperbolic GRU
    (0.2, 0.5) $\begin{matrix}\texttt{eGRU-60-s50}\\ (11584)\end{matrix}$ $\begin{matrix}\texttt{hGRU-55-s50}\\ (9794)\end{matrix}$
    (0.5, 0.2) $\begin{matrix}\texttt{eGRU-60-s50}\\ (11584)\end{matrix}$ $\begin{matrix}\texttt{hGRU-57-s50}\\ (10492)\end{matrix}$
    Table 11: The variants of Euclidean and hyperbolic GRU used as NQS ansatzes to run the second set of VMC experiments involving the 1D $J_{1}J_{2}J_{3}$ Hamiltonian Eq.(55) with $N=30$, $J_{1}=1$ and $(J_{2},J_{3})=(0.2,0.5),(0.5,0.2)$. Only one variant of Euclidean GRU, eGRU-60-s50 with $d_{h}=60$ was used. Three variants of hyperbolic GRU with $d_{h}=53,55,57$ were used in the experiment, but we only list the best performing variant in this Table. The number of parameters of each NQS is listed in brackets under their names.
In the first set of VMC experiments, we only use one variant each for either Euclidean (eGRU-50-s50 with $d_{h}=50$) or hyperbolic GRU (hGRU-50-s50 with $d_{h}=50$) and fix the number of training epochs to 280 in all cases. In the second set of VMC experiments, we explore the performances of hyperbolic GRU variants with $d_{h}=53,55,57$ to choose the best one among them to compare with the Euclidean GRU benchmark, eGRU-60-s50 with $d_{h}=60$, while extending the number of training epochs to 350 for $(J_{2},J_{3})=(0.2,0.5)$ and 500 for $(J_{2},J_{3})=(0.5,0.2)$. Similar to previous cases, in each experiment, hyperbolic GRU can take up to around 5 times longer to complete the training (on the same hardware) than Euclidean GRU with the same $d_{h}$171717while hyperbolic GRU with smaller $d_{h}$ also takes longer to train, up to around 2-3 times, than Euclidean GRU with larger $d_{h}$. All other training details (optimizer settings, saving-of-best-model-only method) remain the same as in the 1D $J_{1}J_{2}$ case and will not be repeated here.
@@ 5.4.2 Results
  • For the first set of VMC experiments involving $(J_{2},J_{3})$ = (0.0, 0.5), (0.2, 0.2), (0.2, 0.5), (0.5, 0.2) with the NQS ansatzes listed in Table 10, the training and inference results are listed in Table 12 and Table 13, respectively. Fig.9 graphically shows the data in both these Tables and offers a comparison between the performances of the 1D Euclidean and 1D hyperbolic GRU. The convergence curve plots for the mean energy and energy variance recorded during the training process are shown in Fig.17 and Fig.18, respectively, in Appendix A.4.
@@
    Table 12: VMC training results (from the first set of experiments with the NQS ansatzes listed in Table 10) for the 1D $J_{1}J_{2}J_{3}$ system with the Hamiltonian Eq.(55) with $N=30$, $J_{1}=1$, and $(J_{2},J_{3})=(0.0,0.2),(0.2,0.2),(0.2,0.5),(0.5,0.2)$. These results are recorded directly from the training and correspond to the parameters of the best saved model. For each ansatz, the mean $E$ is listed first, together with the variance (in brackets). The best result at each $(J_{2},J_{3})$ case is noted in bold.
@@
    Table 13: VMC inference results (from the first set of experiments with the NQS ansatzes listed in Table 10) for the 1D $J_{1}J_{2}J_{3}$ system with the Hamiltonian Eq.(55) with $N=30$, $J_{1}=1$, and $(J_{2},J_{3})=(0.0,0.2),(0.2,0.2),(0.2,0.5),(0.5,0.2)$. For each ansatz, the mean $E$ is listed first, together with the variance (in brackets). The best result at each $(J_{2},J_{3})$ case is noted in bold.
@@
    Figure 9: Comparisons of the performances of Euclidean and hyperbolic GRU NQS ansatzes listed in Table 10 from VMC runs of 1D $J_{1}J_{2}J_{3}$ model with different $(J_{2},J_{3})$ values - clockwise from top: (0.0, 0.5), (0.2, 0.2), (0.5, 0.2), (0.2, 0.5) - for $N=30$ spins. In each subfigure, the mean energy and standard error from training and inference for each NQS ansatz is shown as a dot with an error bar. As is the $J_{1}J_{2}$ case, we do not specify the exact variants (listed in Table 10) in this plot, but choosing instead to label them as ‘Euclidean GRU’ and ‘Hyperbolic GRU’ to emphasize the geometry of the NQS ansatz.
    From Table 12, Table 13 and Fig.9, we note that in both training and inference, hyperbolic GRU definitively outperformed Euclidean GRU (by a large margin) in all 4 out of 4 cases. The energy convergence curves (see Fig.17) corresponding to the hyperbolic GRU ansatz reach convergence at values well below that reached by the curves corresponding to the Euclidean GRU ansatz, while the variance convergence curves of hyperbolic GRU (see Fig.18) show smaller ranges of fluctuations than Euclidean GRU’s.
  • For the second set of VMC experiments involving only ($J_{2},J_{3})=(0.2,0.5),(0.5,0.2)$ with the NQS ansatzes listed in Table 11, the training and inference results are included in Table 14 and Table 15. The comparison of the performances of hyperbolic and Euclidean GRU ansatzes is shown in Fig.10. The convergence curve plots of the second set of VMC experiments for the mean energy and energy variance recorded during the training process are shown in Fig.19 and Fig.20, respectively, in Appendix A.4.
    Ansatz $(J_{2},J_{3})=(0.2,0.5)$ $(J_{2},J_{3})=(0.5,0.2)$
    Euclidean GRU -14.8879 -11.4706
    (1.135) (0.113)
    Hyperbolic GRU -14.6645 -11.5146
    (1.344) (2.536)
    Exact (DMRG) -14.6408 -11.5287
    Table 14: VMC training results (from the second set of experiments with the NQS ansatzes listed in Table 11) for the 1D $J_{1}J_{2}J_{3}$ system with the Hamiltonian Eq.(55) with $N=30$, $J_{1}=1$, and $(J_{2},J_{3})=(0.2,0.5),(0.5,0.2)$. These results are recorded directly from the training and correspond to the parameters of the best saved model. For each ansatz, the mean $E$ is listed first, together with the variance (in brackets). The best result at each $(J_{2},J_{3})$ case is noted in bold.
    Ansatz $(J_{2},J_{3})=(0.2,0.5)$ $(J_{2},J_{3})=(0.5,0.2)$
    Euclidean GRU -14.6828 -11.2061
    (1.000) (0.403)
    Hyperbolic GRU -14.4243 -11.3070
    (0.640) (0.065)
    Exact (DMRG) -14.6408 -11.5287
    Table 15: VMC inference results (from the second set of experiments with the NQS ansatzes listed in Table 11) for the 1D $J_{1}J_{2}J_{3}$ system with the Hamiltonian Eq.(55) with $N=30$, $J_{1}=1$, and $(J_{2},J_{3})=(0.2,0.5),(0.5,0.2)$. For each ansatz, the mean $E$ is listed first, together with the variance (in brackets). The best result at each $(J_{2},J_{3})$ case is noted in bold.
    Figure 10: Comparisons of the performances of Euclidean and hyperbolic GRU NQS ansatzes listed in Table 11 from VMC runs of 1D $J_{1}J_{2}J_{3}$ model with different $(J_{2},J_{3})$ values - from left to right: (0.2, 0.5), (0.5, 0.2) - for $N=30$ spins. In each subfigure, the mean energy and standard error from training and inference for each NQS ansatz is shown as a dot with an error bar. As before, we do not specify the exact variants (listed in Table 11) in this plot, but choosing instead to label them as ‘Euclidean GRU’ and ‘Hyperbolic GRU’ to emphasize the geometry of the NQS ansatz.
    From Table 14, Table 15 and Fig.10, for $(J_{2},J_{3})$ = (0.2, 0.5), the hyperbolic GRU ansatz hGRU-55-s50 outperforms the Euclidean GRU ansatz eGRU-60-s50 in the training process, as is evidence from the closer mean energy value obtained during training, as well as from the energy convergence curve shown in the first subfigure of Fig.19, which shows the curve corresponding to hGRU-55-s50 reaching convergence faster with a narrower fluctuation range than the curve corresponding to eGRU-60-s50. In this case, it must be noted that the Euclidean GRU ansatz used, eGRU-60-s50, has around 20 percent more parameters (11584 versus 9794) than the hyperbolic GRU ansatz, hGRU-55-s50. For $(J_{2},J_{3})$ = (0.5, 0.2), the hyperbolic GRU ansatz hGRU-57-s50 outperforms the Euclidean GRU ansatz eGRU-60-s50 in both the training and inference processes. In this case, hyperbolic GRU has 10 percent fewer parameters than Euclidean GRU (10492 versus 11584).
@@
From the two sets of VMC experiments for the 1D $J_{1}J_{2}J_{3}$ model, we observe that the obtained results showing that hyperbolic GRU almost always outperform Euclidean GRU are in agreement with the ones in the 1D $J_{1}J_{2}$ case in which hyperbolic GRU outperfomed its Euclidean version in all cases when $J_{2}\neq 0$, i.e. when the second neighbor interaction is nonvanishing and the Hamiltonian exhibits a hierarchy in the interaction structure comprised of the first and second neighbor interactions. In all four $J_{1}J_{2}J_{3}$ settings where there exist an even more pronounced hierarchy in the interaction structure of the Hamiltonian that comprises the first, second and third nearest neighbor interactions, the clear outperformance of the hyperbolic GRU ansatz reinforces the hypothesis that the hyperbolic geometry of this neural network plays a determining role in handling the type of Hamiltonians with hierarchical structures such as the 1D $J_{1}J_{2}$, $J_{1}J_{2}J_{3}$ systems. As already noted in the 1D $J_{1}J_{2}$ case, an interesting task would be to verify whether hyperbolic GRU carries on outperforming Euclidean GRU for the 2D $J_{1}J_{2}J_{3}$ models on different lattice geometries, but this will require the generalization of the 1D hyperbolic GRU network to a 2D version, similar to the generalization of the 1D Euclidean RNN cell to the 2D version in Eq.52 as done in [9].